OpenStudy (anonymous):

A teacher gave her class two exams; 60% of the class passed the second exam, but only 48% of the class passed both exams. What percent of those who passed the second exam also passed the first exam?

OpenStudy (anonymous):

here is an easy trick to know since you don't know how many students were in the class, pick a number and work with the number a good number to pick is 100

OpenStudy (anonymous):

48 students passed both exams, 60 past the first one

OpenStudy (anonymous):

can you help me on what to do next

OpenStudy (anonymous):

i made a typo there, should have been 48 students passed both exams, 60 passed the second one out of the 60 students who passed the second exam, 48 also passed the first one you are asked What percent of those who passed the second exam also passed the first exam?

OpenStudy (anonymous):

you job it to find the percent for \(\frac{48}{60}\)

OpenStudy (anonymous):

divide and move the decimal over two places

OpenStudy (anonymous):

or you can say \[\frac{48}{60}=\frac{8}{10}=.8=80\%\]

OpenStudy (anonymous):

clear enough?

OpenStudy (anonymous):

yea thank you

OpenStudy (anonymous):

yw